# Action, entropy and pair creation rate of charged black holes in de Sitter space

### v2

@@ 1 Introduction
Here $\mathcal{A}_{+}$ is the area of the (outer) black hole horizon and $\mathcal{A}_{++}$ is the cosmological horizon area. This formula suggests that the total entropy of the RNdS static patch between the two horizons is given by the sum of the horizon entropies. Applying this result we then compute the pair production rate of any static black hole in de Sitter space using the formalism of constrained instantons in section 4. We end with a comment on the Festina Lente bound FestinaLente:2019 in section 5, and a short summary and open questions in section 6.
@@ 3.1 Hartle-Hawking wavefunction and instantons
We are interested in computing the pair creation rate of charged black holes in a de Sitter background for any mass and charge. The pair creation of such black holes is described by the propagation from nothing to a spatial surface $\Sigma$ with topology $S^{D-2}\times S^{1}$ for generic mass and charge (and $\Sigma$ has topology $S^{D-2}\times R$ for extreme and ultracold black holes). Given certain boundary conditions on $\Sigma$ fixing the mass and charge, the probability amplitude for this process is given by the no-boundary Hartle-Hawking wavefunction, which can be formally represented as a Euclidean path integral Hawking:1995ap ; Mann:1995vb
@@ 3.2 On-shell Euclidean action
We emphasize that this expression is independent of the Euclidean time period $\beta$, and that it holds in any number of dimensions for arbitrary mass and charge. This is the action in the canonical ensemble where the temperature and charge are kept fixed. [^3] In the grand canonical ensemble that fixes the electric potential instead of the charge, there is no boundary term for the gauge field and the on-shell action of $\mathcal{M}_{1/2}$ is given by (39) instead. The same result for the on-shell action $I_{1/2,RNdS}$ was obtained in four dimensions in Chao:1997osu ; Wang:2022sbp , which was used to compute the pair creation of fixed mass and charge black holes in de Sitter. For zero charge the result (43) is consistent with earlier findings for the action of Euclidean Schwarzschild-de Sitter space Chao:1997em ; Bousso:1998na ; Gregory:2014 ; SdSaction . Further, the on-shell action of the extreme Euclidean solutions (proper instantons) was computed before in Mellor:1989wc ; Hawking:1995ap ; Mann:1995vb ; Cardoso:2004uz , which we discuss in more detail below.
@@ 3.3 Extreme cases: cold, lukewarm, Nariai and ultracold instantons
In the literature the action is usually only evaluated for the extreme Euclidean RNdS solutions: cold, lukewarm, Nariai and ultracold instantons. This is because they are proper stationary points of the Euclidean Einstein action, i.e., they are regular and hence satisfy the Einstein equation everywhere. Below we specialize to these true instanton cases and compare our results for the on-shell action to those in the literature, especially to the results in the comprehensive paper by Mann and Ross Mann:1995vb . From our general expression (43) it follows that the on-shell actions of the extremal, lukewarm, Nariai and ultracold instantons are, respectively,
@@
where $I_{NN}$ denotes the action of the neutral Nariai solution ($\rho_{NN}=0$) and $I_{LN}$ the action of the lukewarm-Nariai solution ($\rho_{LN}=\frac{\sqrt{3}}{2}$). In Figure 6 we plot the fraction of the proper instanton actions and the de Sitter action against the charge parameter $\rho=q/q_{U}$. We note the actions of the lukewarm (red line) and cold (green line) solutions coincide at $\rho_{dS}=0$ and are equal to the de Sitter action $I_{dS}$ in that case. Further, the actions of the lukewarm and Nariai (blue line) solutions coincide for the lukewarm-Nariai solution (yellow point). Moreover, we can see from the figure that the actions of the extremal and Nariai solutions are the same at $\rho=1$, which corresponds to the ultracold solution (magenta point). The extremal and Nariai solutions also have the same action at $\rho=\frac{\sqrt{5}}{3}$, but this is not a special point in the shark fin diagram. Figure 6 can be contrasted with Figure 3 in a paper by Bousso and Hawking Bousso:1996au where the fraction $I/I_{dS}$ is plotted against the square of the charge parameter (in fact, the dimensionless quantity $q^{2}\Lambda).$ In that figure made by Bousso and Hawking (which is based on the computations of the action by Mann and Ross Mann:1995vb ) the Nariai branch and the extremal branch do not meet at the ultracold point, which seems at odds with the shark fin diagram.
@@ 4 Constrained instantons and the pair creation rate
$\delta_{g}I_{E}[g,A]+\lambda_{m}\,\delta_{g}{\cal C}[g,A]+\lambda_{q}\,\delta_{g}{\cal D}[g,A]=0\,,$
$\delta_{A}I_{E}[g,A]+\lambda_{m}\,\delta_{A}{\cal C}[g,A]+\lambda_{q}\,\delta_{A}{\cal D}[g,A]=0\,,$
@@ 5 A comment on the Festina Lente bound
As explained in Aalsma:2018 , using an s-wave tunneling formalism in flat and AdS space, this is what one indeed finds, even in the so-called “superradiant” regime at low energies, where $\Delta S>0$, describing a fast, entropically favourable, decay of the charge of the black hole. We expect an analogous line of reasoning to hold just as well in de Sitter space. Indeed, in FestinaLente:2019 one concludes, supported by explicit calculations of the Schwinger pair creation rate, that the adiabatic, small mass regime should be avoided, [^5] This should coincide with the superradiant regime, as explained in Aalsma:2018 . otherwise the decay to singular states to the right of the Nariai boundary would be possible, violating the Cosmic Censorship Conjecture and prohibiting an effective description of the (slow) decay. This sets a lower “Festina Lente” bound on the mass of an elementary excitation, in contrast to the Weak Gravity Conjecture, which sets an upper bound on the mass of an elementary excitation.
This supports the conjecture that an s-wave tunneling calculation in de Sitter space should be able to reproduce the effective evolution inside the allowed shark fin region, as long as one excludes the superradiant regime by requiring the absence charged particles that are too light. In other words, we would like to suggest that the Festina Lente bound can be derived (and understood) more precisely by identifying the microscopic mass and charge parameters for which $\Delta S$ vanishes. This would be interesting to check explicitly, especially in the ultracold limit, where the macroscopic effective decay, described in terms of $dq/dm$, is fixed uniquely and equals
@@ 6 Summary and conclusions
Finally, with the result for the Euclidean action, in arbitrary dimensions, and the expectation that thermodynamics should be governing the decay of charged black holes in de Sitter, we argued that a tunneling calculation similar to Aalsma:2018 should be able to explicitly reproduce the slow effective decay of charged de Sitter black holes within the allowed shark fin region as long as one avoids the superradiant regime. This would be particularly interesting to check in the ultracold limit, where the effective evolution describing the decay is fixed uniquely. We conjecture that by using the tunneling formalism it should be possible to derive a more precise version of the Festina Lente type bound in arbitrary dimensions.
@@ Appendix B Arbitrary dimensions
Next, by solving $f(r_{+})=f(r_{++})=0$, the mass and charge parameters can be expressions in terms of the outer horizon radius $r_{+}$, the cosmological horizon radius $r_{++}$ and de Sitter radius $l$
@@
$\left(\frac{m_{LN}}{m_{U}}\right)^{2}=3^{D-3}(2D-5)^{2}\frac{(D-2)^{D-3}}{(3D-8)^{D-1}}\,,\qquad\left(\frac{q_{LN}}{q_{U}}\right)^{2}=3^{D-3}\left(\frac{D-2}{3D-8}\right)^{D-2}\,.$ (88)
Finally, the horizon radius of the lukewarm-Nariai black hole is given by

### v3

@@ 1 Introduction
Here $\mathcal{A}_{+}$ is the area of the (outer) black hole horizon and $\mathcal{A}_{++}$ is the cosmological horizon area. This formula suggests that the total entropy of the RNdS static patch between the two horizons is given by the sum of the horizon entropies. Applying this result we then compute the pair production rate of any static black hole in de Sitter space using the formalism of constrained instantons in section 4. We end with a comment on the Festina Lente bound Montero:2019ekk in section 5, and a short summary and open questions in section 6.
@@ 3.1 Hartle-Hawking wavefunction and instantons
We are interested in computing the pair creation rate of charged black holes in a de Sitter background for any mass and charge. [^3] From a global de Sitter perspective this is naturally interpreted as creating a pair of opposite charge, opposite mass, black holes, due to the opposite time orientation in two conjugate static patches. In a single static patch region this instead corresponds to the spontaneous creation of a single charged black hole. The pair creation of such black holes is described by the propagation from nothing to a spatial surface $\Sigma$ with topology $S^{D-2}\times S^{1}$ for generic mass and charge (and $\Sigma$ has topology $S^{D-2}\times R$ for extreme and ultracold black holes). Given certain boundary conditions on $\Sigma$ fixing the mass and charge, the probability amplitude for this process is given by the no-boundary Hartle-Hawking wavefunction, which can be formally represented as a Euclidean path integral Hawking:1995ap ; Mann:1995vb
@@ 3.2 On-shell Euclidean action
We emphasize that this expression is independent of the Euclidean time period $\beta$, and that it holds in any number of dimensions for arbitrary mass and charge. This is the action in the canonical ensemble where the temperature and charge are kept fixed. [^4] In the grand canonical ensemble that fixes the electric potential instead of the charge, there is no boundary term for the gauge field and the on-shell action of $\mathcal{M}_{1/2}$ is given by (39) instead. The same result for the on-shell action $I_{1/2,RNdS}$ was obtained in four dimensions in Chao:1997osu ; Wang:2022sbp , which was used to compute the pair creation of fixed mass and charge black holes in de Sitter. We expect the creation of multiple charged black hole in de Sitter to be further suppressed with (a multiple power of) the same exponential factor, due to the fact that the exponential de Sitter expansion will create multiple independent static patches that contain just a single charged black hole. This expectation is in fact supported by the explicit construction of the (dynamical) solution for multiple de Sitter charged black holes Kastor:1992nn , generalizing the Majumdar-Papapetrou solutions in flat space. For zero charge the result (43) is consistent with earlier findings for the action of Euclidean Schwarzschild-de Sitter space Chao:1997em ; Bousso:1998na ; Gregory:2014 ; SdSaction . Further, the on-shell action of the extreme Euclidean solutions (proper instantons) was computed before in Mellor:1989wc ; Hawking:1995ap ; Mann:1995vb ; Cardoso:2004uz , which we discuss in more detail below.
@@ 3.3 Extreme cases: cold, lukewarm, Nariai and ultracold instantons
In the literature the action is usually only evaluated for the extreme Euclidean RNdS solutions: cold, lukewarm, Nariai and ultracold instantons. This is because they are proper stationary points of the Euclidean Einstein action, i.e., they are regular and hence satisfy the Einstein equation everywhere. Below we specialize to these true instanton cases and compare our results for the on-shell action to those in the literature, especially the comprehensive paper by Mann and Ross Mann:1995vb . From our general expression (43) it follows that the on-shell actions of the extremal, lukewarm, Nariai and ultracold instantons are,
@@
where $I_{NN}$ denotes the action of the neutral Nariai solution ($\rho_{NN}=0$) and $I_{LN}$ the action of the lukewarm-Nariai solution ($\rho_{LN}=\frac{\sqrt{3}}{2}$). In Figure 6 we plot the fraction of the proper instanton actions and the de Sitter action against the charge parameter $\rho=q/q_{U}$. We note the actions of the lukewarm (red line) and cold (green line) solutions coincide at $\rho_{dS}=0$ and are equal to the de Sitter action $I_{dS}$ in that case. Further, the actions of the lukewarm and Nariai (blue line) solutions coincide for the lukewarm-Nariai solution (yellow point). Moreover, we can see from the figure that the actions of the extremal and Nariai solutions are the same at $\rho=1$, which corresponds to the ultracold solution (magenta point). The extremal and Nariai solutions also have the same action at $\rho=\frac{\sqrt{5}}{3}$, but this is not a special point in the shark fin diagram. Figure 6 can be contrasted with Figure 3 in a paper by Bousso and Hawking Bousso:1996au where the fraction $I/I_{dS}$ is plotted against the square of the charge parameter (in fact, the dimensionless quantity $q^{2}\Lambda).$ In that figure by Bousso and Hawking (which is based on the computations of the action by Mann and Ross Mann:1995vb ) the Nariai branch and the extremal branch do not meet at the ultracold point, which seems at odds with the shark fin diagram.
@@ 4 Constrained instantons and the pair creation rate
$\delta_{g}I[g,A]+\lambda_{m}\,\delta_{g}{\cal C}[g,A]+\lambda_{q}\,\delta_{g}{\cal D}[g,A]=0\,,$
$\delta_{A}I[g,A]+\lambda_{m}\,\delta_{A}{\cal C}[g,A]+\lambda_{q}\,\delta_{A}{\cal D}[g,A]=0\,,$
@@ 5 A comment on the Festina Lente bound
As explained in Aalsma:2018 , using an s-wave tunneling formalism in flat and AdS space, this is what one indeed finds, even in the so-called “superradiant regime” at low energies, where $\Delta S>0$, describing a fast, entropically favourable, decay of the charge of the black hole. We expect an analogous line of reasoning to hold just as well in de Sitter space. Indeed, in Montero:2019ekk one concludes, supported by explicit calculations of the Schwinger pair creation rate, that the adiabatic, small mass regime should be avoided, [^6] This should coincide with the superradiant regime, as explained in Aalsma:2018 . As for rotating black holes, the superradiant regime for charged black holes allows energy to be extracted by scattering a beam of charged particles off the charged black hole, through stimulated pair creation in a region close to the black hole. This charged black hole ‘generalized ergosphere’ region will depend on the properties of the charged particle flux, as opposed to the universal ergosphere of a rotating black hole Denardo:1973pyo . otherwise the decay to singular states to the right of the Nariai boundary would be possible, violating the Cosmic Censorship Conjecture and prohibiting an effective description of the (slow) decay. This sets a lower “Festina Lente” bound on the mass of an elementary excitation, in contrast to the Weak Gravity Conjecture, which sets an upper bound on the mass of an elementary excitation.
This supports the conjecture that an s-wave tunneling calculation in de Sitter space should be able to reproduce the effective evolution inside the allowed shark fin region, as long as one excludes the superradiant regime by requiring the absence of charged particles that are too light. In other words, we would like to suggest that the Festina Lente bound can be derived (and understood) more precisely by identifying the microscopic mass and charge parameters for which $\Delta S$ vanishes. This would be interesting to check explicitly, especially in the ultracold limit, where the macroscopic effective decay, described in terms of $dq/dm$, is fixed uniquely and equals
@@ 6 Summary and conclusions
Finally, with the result for the Euclidean action, in arbitrary dimensions, and the expectation that thermodynamics should be governing the decay of charged black holes in de Sitter, we argued that a tunneling calculation similar to Aalsma:2018 should be able to explicitly reproduce the slow effective decay of charged de Sitter black holes within the allowed shark fin region as long as one avoids the superradiant regime. This would be particularly interesting to check in the ultracold limit, where the effective evolution describing the decay is fixed uniquely. We conjecture that by using the tunneling formalism it should be possible to derive a more precise version of the Festina Lente type bound in arbitrary dimensions (see our follow-up paper Aalsma:2023mkz ).
@@ Appendix B Arbitrary dimensions
Next, by solving $f(r_{+})=f(r_{++})=0$, the mass and charge parameters can be expressed in terms of the outer horizon radius $r_{+}$, the cosmological horizon radius $r_{++}$ and de Sitter radius $l$
@@
$\left(\frac{m_{LN}}{m_{U}}\right)^{2}=3^{D-3}(2D-5)^{2}\frac{(D-2)^{D-3}}{(3D-8)^{D-1}}\,,\qquad\left(\frac{q_{LN}}{q_{U}}\right)^{2}=3^{D-3}\left(\frac{D-2}{3D-8}\right)^{D-2}\,,$ (86)
such that the horizon radius of the lukewarm-Nariai black hole is given by
@@
a) Shark fins
b) Normalized shark fins
Figure 10: Shark fin phase diagrams for $D=4$ (green), $D=5$ (turquoise), $D=6$ (blue) and $D=7$ (purple) obtained with (7). In the limit $D\rightarrow\infty$, the shark fins in the left panel get smaller and smaller, whereas the ones in the right panel converge to the values in (88) and (89).
Finally, defining $\nu_{N},\nu_{E}\in[0,1]$ via $r_{N}\equiv\nu_{N}\,r_{U}+(1-\nu_{N})\,r_{N}^{0}$ and $r_{E}\equiv\nu_{E}^{1/D}\,r_{U}$, where $r_{N}^{0}=\sqrt{\frac{D-3}{D-1}}l$ stands for the neutral Nariai radius and the radii $r_{E,N}$ are related to $m_{E,N}$ and $q_{E,N}$ by (7), allows to obtain the limits of the ratios
$\frac{m_{N}}{m_{U}}$ $\xrightarrow{D\rightarrow\infty}\frac{1}{2}(1+\nu_{N})\,e^{\frac{1}{2}(1-\nu_{N})}\,,\qquad\frac{q_{N}}{q_{U}}\xrightarrow{D\rightarrow\infty}\sqrt{\nu_{N}}\,e^{\frac{1}{2}(1-\nu_{N})}$ (88)
$\frac{m_{E}}{m_{U}}$ $\xrightarrow{D\rightarrow\infty}\nu_{E}(1-\ln\nu_{E})\,,\qquad\qquad\,\,\,\frac{q_{E}}{q_{U}}\xrightarrow{D\rightarrow\infty}\nu_{E}\sqrt{1-2\ln\nu_{E}}\,.$ (89)
These values set the boundaries of the normalized shark fin phase diagram in Figure 10 in the $D\rightarrow\infty$ limit.